OpenStudy (goformit100):

A survey that 63% of the American like cheese whereas 76% like apples. If x% of the Americans like both cheese and apples, find the value of x.

OpenStudy (goformit100):

@mary.rojas

OpenStudy (mary.rojas):

let us take % as no. of people. Now no.of americans who like cheese , n(c) =63 no.of americans who like apples ,n(a)=76 and total no. of people , n(cUa)=100 (since 100%) you need to find the intersection of the sets cheese and apples which is equal to n(a)+n(c)-n(cUa)=63+76-100=39 hence X=39 Posted by saiarivalan_4...(student), on 5/11/12 http://www.meritnation.com/ask-answer/question/a-survey-shows-that-63-of-americans-like-cheese-whereas/sets/3235480

OpenStudy (anonymous):

thanks for citing

OpenStudy (mary.rojas):

lol your welcome. Am not a plagiarizer :)